i. A 150MLD water treatment plant needs a dosage of 2 ppm chlorine to maintain a residual chlorine of 1.2 mg/L. What is the quantity of chlorine required (in kg) for the whole month of August 2020?

ii. Describe three (3) approaches that can be used for water disinfection.
Answer #1

I) Given,

Q = 150 MLD

Dosage = 2 ppm = 2 mg/L

1 ppm = 1 mg/L

Residual chlorine = 1.2 mg/L

1 MLD * 1 mg/L = 1 kg/day

Quantity of chlorine per one day = Q * Chlorine Demand = Q * (Dosage - Residual Chlorine) = 150 * (2.0 - 1.2) = 120 kg/day

Quantity of chlorine required for the whole month of August 2020 = 120 * 31 = 3720 kg